OpenStudy (anonymous):

What is the slope of the line passing through the point (4, −2) and perpendicular to the line y = 4x + 2? Write your answer as a fraction.

OpenStudy (anonymous):

-1/2

OpenStudy (anonymous):

a line perpendicular to y=4x+2 has the form y=(-1/4)x+b so solve for b using the point -2=(-1/4)4+b -2=-1+b b=-1 so the line is y=(-1/4)x-1 the slope of the line is then -1/4